• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kelleher

    Log entry time 23:59:28 on April 15, 2007

    Entry number 199199

    keyword=Swing Shift Summary

    Sunday, April 15 Swing Shift

    16:00 Shift began with no beam
    16:50 MCC called, and they are ready to send beam to us
    16:55 Had to (re)replace CAMAC power supply, taking beam away.
    17:10 MCC called, and they are ready to send beam to us
    17:15 Spot checked at low current (10uA), asked for 95uA
    20:55 reduced current (heater load was low), then opened JT valve and
    came back to 95uA
    23:15 ROC4 hung up, rebooted, killed CODA, restarted CODA

    Very nice quiet running. Apart from Q1 and CODA, no alarms and no surprises.

    Run Summary

    Run Events Notes
    -----------------------------
    2510 3k Spot++ check
    2511 1.4M 95uA production run (did short replay (replay.C))
    2512 1.4M 95uA production run
    2513 1.4M 95uA production run
    2514 1.4M 95uA production run
    2515 1.4M 95uA production run (did short replay (replay.C))
    2516 1.4M 95uA production run
    2517 1.4M 95uA production run
    2518 300k 95uA production run (Left Q1 trip)
    2519 1.4M 95uA production run
    2520 1.4M 95uA production run
    2521 400k 95uA production run (did short replay (replay.C)) CODA crashed.
    2522 1.4M 95uA production run
    2523 95uA production run (started at 23:55)